# AMPLIFY:Attention-based Mixup for Performance Improvement and Label Smoothing in Transformer

### v2

@@ Abstract.
Mixup is an effective data augmentation method that generates new augmented samples by aggregating linear combinations of different original samples. However, if there are noises or aberrant features in the original samples, Mixup may propagate them to the augmented samples, leading to over-sensitivity of the model to these outliers 5.3. To solve this problem, this paper proposes a new Mixup method called AMPLIFY. This method uses the Attention mechanism of Transformer itself to reduce the influence of noises and aberrant values in the original samples on the prediction results, without increasing additional trainable parameters, and the computational cost is very low, thereby avoiding the problem of high resource consumption in common Mixup methods such as Sentence Mixup (Guo et al., 2019a). The experimental results show that, under a smaller computational resource cost, AMPLIFY outperforms other Mixup methods in text classification tasks on 7 benchmark datasets, providing new ideas and new ways to further improve the performance of pre-trained models based on the Attention mechanism, such as BERT, ALBERT, RoBERTa, and GPT. Our code can be obtained at https://github.com/kiwi-lilo/AMPLIFY.
Mixup, Attention Mechanism, Data Augmentation, Model Robustness, Label Smoothing
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXccs: Computing methodologies Natural language processingccs: Computing methodologies Language resourcesccs: Computing methodologies Modeling methodologies
1. Introduction
@@
Data augmentation techniques are widely used in modern machine learning field to improve the predictive performance and robustness of computer vision and natural language processing (NLP) models by adding feature noise to the original samples. Compared to computer vision tasks, NLP tasks face samples with more complex data structures, semantic features, and semantic correlations, as natural language is the main channel of human communication and reflection of human thinking. (Li et al., 2022) Therefore, NLP models are more sensitive to the quality of the dataset, and often require a variety of data augmentation techniques to improve the model’s generalization ability, adaptability, and robustness to new data in practical engineering applications. Common text data augmentation techniques include random synonym replacement  (Zhang et al., 2015), back-translation  (Xie et al., 2020), random word insertion and deletion, etc.  (Wei and Zou, 2019). The core idea of these methods is to generate more training samples by performing a series of feature transformations on the original samples, allowing the model to learn and adapt to changing feature information, and enabling the model to better deal with and process uncertainty in the samples. Additionally, by selectively increasing the size of the training set, data augmentation techniques can effectively alleviate the negative impact of limited data and imbalanced data classification on model prediction performance.
@@
Standard Mixup is a simple and effective image augmentation method first proposed in 2017  (Zhang et al., 2017). It aggregates two images and their corresponding labels by linear interpolation to generate a new augmented image and its pseudo-label. The main advantage of the standard Mixup technique is that it purposefully generates beneficial noise by using the weighted average of features in the original samples. After adapting to this noise, the model becomes less sensitive to other noise in the training samples, thereby improving the model’s generalization ability and robustness. Additionally, because it can generate more augmented samples by aggregating different original sample pairs, even at a high data augmentation magnitude, there is no duplicate aggregation result. This greatly enhances the diversity of augmented samples, thereby effectively improving the training efficiency of the model and reducing the risk of overfitting.In the NLP field, the standard Mixup technique can be broadly divided into two categories: Input Mixup  (Yun et al., 2019) and Hidden Mixup  (Verma et al., 2019). Input Mixup increases the number of samples in the training set input to the model. The specific process involves padding all text sequences in the training dataset to the same length. After completing word embedding processing, random pairs of sequences are combined. Mixup operations are then applied to the two vectorized samples in each sequence pair, along with their corresponding classification labels, generating augmented samples and their pseudo-labels. Finally, the augmented samples and pseudo-labels are fed into the model for training. Input Mixup performs Mixup operations during the word embedding stage of text preprocessing and only involves semantic features in the word vector space. Its main objective is to increase the training set size and reduce overfitting. Hidden Mixup applies Mixup operations in the hidden layers of the model. The specific process involves randomly selecting two sequences, $x$ and $x^{\prime}$, from the training set. After processing through the model, a random layer is selected from the model’s hidden layers, and the output features, $h$ and $h^{\prime}$, of $x$ and $x^{\prime}$ at that layer are extracted. Mixup is then performed on $h$, $h^{\prime}$, and their labels, resulting in linearly interpolated mixed features and corresponding pseudo-labels.
@@
Finally, the mixed features are fed into subsequent hidden layers. Hidden Mixup performs Mixup during the feature extraction stage of text processing and only involves specific hidden layers in the model. Its main objective is to enhance the model’s generalization ability without requiring additional input data. In summary, Input Mixup increases the sample quantity, while Hidden Mixup increases feature diversity. Both methods focus on improving model performance through Mixup operations, albeit with different emphases.  (Verma et al., 2019)’s research has shown that performing Mixup operations at the input layer of the model leads to underfitting, while performing Mixup at deeper layers makes the training easier to converge. However, when the original samples contain noise or outliers, performing Mixup may result in generated augmented samples containing more regular noise or outliers due to the principle of linear interpolation. If the model learns too much from this noise and outliers, it may cause obvious overfitting and prediction bias, thereby reducing the model’s generalization performance. The basic idea of Standard Mixup is to linearly aggregate two samples $\left(x_{i},y_{i}\right)$ and $\left(x_{j},y_{j}\right)$ in the training set $\mathbb{D}_{train}$, to generate a new sample $\left(x_{\mathrm{mix}},y_{\mathrm{mix}}\right)$ as an augmented input to train the network, where $x$ represents the sample and $y$ represents its corresponding label. This aggregation process can be expressed by the following formulae:
@@
Figure 1. A schematic of AMPLIFY. In each encoder block of the Transformer, the forward-propagated input data is duplicated and re-ordered according to the label-mixing order after obtaining the results of the Muti-Head Attention, and then Mixup operation is performed. No changes are made to other network structures. Similarly, each decoder block can also perform the same operation.
@@
Multi-Head Attention (MHA) (Vaswani et al., 2017) is a technique commonly used in sequence-to-sequence models to calculate feature correlations. Namely, it can calculate the attention weights for each position in the sequence in parallel using different attention heads, and then obtain the representation of the entire sequence by summing the weighted values. Specifically, MHA helps the model establish correspondences between features in the input and output sequences, thereby improving the expressiveness and accuracy of the model. It can also attend to different features at different positions in the input sequence without taking into account their distance, assigning different weights to different parts of the input to capture information relevant to the output sequence. To address the problem of noise and out-of-distribution values in original samples, which affect the aggregated samples in standard Mixup, our AMPLIFY method uses the Hidden Mixup idea to perform Mixup operations on the hidden layers of the Transformer block. Since the output of the MHA already includes the model’s attention to different parts of the input sequence, it can guide the model on which features to retain and which to discard when aggregating different samples, ultimately integrating more critical features into the generated output sequence. Therefore, AMPLIFY duplicates the MHA output results of the sample sequence in the same batch, shuffles the order of these copies, and then performs Mixup on them with the original output results. This allows the model to aggregate the correlations and attention of two sample features multiple times at different levels, thereby obtaining more reasonable weights for each position in the aggregated sequence and avoiding unnecessary noise or feature information loss during the Mixup process. Additionally, since the Attention mechanism itself is a method of weighting the input features, AMPLIFY is more natural and effective intuitively.
@@
Since AMPLIFY aggregates the features of the augmented sample and the corresponding original sample in the hidden layers of the model, it has lower computational cost and fewer parameters compared to other data augmentation methods. It also avoids the problem of increased time and resource consumption caused by standard Mixup methods 5. In addition, experimental results have shown that our method can better learn the key features of the input sequence, improving the generalization ability and prediction performance of the model  2. For example, on the MRPC dataset, AMPLIFY’s average accuracy is 1.83% higher than the baseline, and 1.04%, 1.72%, and 1.46% higher than EmbedMix, SentenceMix, and TMix, respectively.
1.1. Motivation
@@ 2. Related Work
Data Augmentation: refers to a class of methods that generate new data with certain semantic relevance based on the features of existing data. By adding more augmented data, the overall prediction performance of the model can be improved and the robustness of the model can be enhanced. In the case of a limited size of the training set, data augmentation techniques are more effective because they can significantly reduce the risk of overfitting and improve the generalization ability of the model. However, the inherent difficulty of NLP tasks makes it difficult to construct data augmentation methods similar to those in the computer vision field (such as cropping and flipping), which may significantly alter the semantics of the text and make it difficult to balance the quality of the data with the diversity of the features. Currently, data augmentation methods in the NLP field can be roughly divided into three categories: Paraphrasing  (Wang and Yang, 2015), Noising  (Wei and Zou, 2019), and Sampling  (Min et al., 2020). Among them, Paraphrasing methods involve certain transformations of the characters, words, phrases, and sentence structures in the text while trying to retain the semantics of the original text. However, such methods may lead to differences in textual semantics in different contexts, for example, substituting "I eat an iPhone every day" for "I eat an apple every day" obviously does not make sense. Noising methods aim to add some continuous or discrete noise to the sentence while keeping the label unchanged. Although this method has little impact on the semantics of the text, it may have a significant impact on the basic structure and grammar of the sentence, and there are also certain limitations in improving the diversity of features. For example, adding noise to the sentence "i like watching sunrise" to turn it into "i like, watching jogging sunrise" would impair the grammatical integrity of the original sentence. The goal of Sampling methods is to select samples based on the feature distribution of the existing data and use these samples to augment new data. This method needs to define different selection strategies manually according to the features of different datasets, so its application range is limited and the diversity of the features obtained by augmentation is relatively poor.
@@
Mixup: is a noise-based data augmentation technique introduced by  (Zhang et al., 2017). It was first introduced in the field of computer vision and is now widely used in the entire field of deep learning. Its purpose is to improve the predictive performance of neural network models by generating regularly noised samples. The standard Mixup method generates a new augmented image by linearly interpolating two existing images and their labels, encouraging the model to learn more generalized decision boundaries. To bring the idea of standard Mixup data augmentation to the NLP field,  (Guo et al., 2019a) modified the underlying operations of standard Mixup and proposed two methods for applying Mixup in NLP tasks: EmbedMix, which performs Mixup in the word embedding space, and SentenceMixup, which performs Mixup after information aggregation in the final hidden layer.  (Sun et al., 2020) first proposed a strategy for doing Mixup on transformer-based pre-trained text models and concluded that the smaller the data volume, the greater the improvement in model performance due to Mixup.  (Chen et al., 2020) proposed a Mixup method for semi-supervised text classification called TMix, which uses two different sets of data to randomly select a layer in all hidden layers of the model and perform Mixup on its features. SeqMix, proposed by  (Zhang et al., 2020b), is a technique for improving feature diversity during active learning by mixing sequences in the feature space of hidden layers. It provides an effective solution for Mixup at the subsequence level and a method for augmenting the labels of active sequences.
@@
Attention: is a commonly used technique in machine learning that allows models to focus on important information in sequence data and learn the underlying features of that information. In the NLP field, Attention assigns a weight to each hidden state and determines which feature state the output should focus on based on the varying weights. This helps the model better understand key features in a sentence, ultimately improving the performance of sequence models. Recently, some computer vision models have attempted to combine Mixup with saliency-based Attention mechanisms, such as attentive-CutMix  (Walawalkar et al., 2020), puzzle-Mix  (Kim et al., 2020), and saliency-CutMix  (Uddin et al., 2020). One successful case of introducing Mixup to the mainstream pre-trained model, Vision Transformer (ViT), is TransMix  (Chen et al., 2022), which dynamically reassigns label weights based on the response of each data point in the attention map. Inputs that are better focused in the attention map are assigned higher values in the mixed labels.
@@
Hidden Mixup: Recent research in data augmentation has explored the use of Mixup in the hidden layers of deep learning models. For example,  (Verma et al., 2019) proposed a Mixup augmentation algorithm called Manifold Mixup, which trains neural networks on interpolated hidden representations and encourages the network to maintain uncertainty about the size of the unobserved feature representation space during training. This causes the feature representations of real training samples to be concentrated in a low-dimensional subspace, resulting in more discriminative aggregated features.  (Yoon et al., 2021) proposed a data augmentation algorithm called SsMix, which performs span-based mixing on input text sequences. This method selects the least salient span in text $A$ and replaces it with the most salient span of the same length in text $B$, while preserving most of the important tokens in both sequences. However, current research has not yet explored the combination of Mixup and MHA in the hidden layer of Transformer architecture.
3. Method
@@
    (1) $\textbf{M}_{B}^{o}:\mathbb{D}_{train}\rightarrow\left\{\left\langle x_{i}^{o},y_{i}^{o}\right\rangle\right\}_{i=1}^{l}=\left\{\left\langle x_{1}^{o},y_{1}^{o}\right\rangle,\left\langle x_{2}^{o},y_{2}^{o}\right\rangle,\ldots,\left\langle x_{l}^{o},y_{l}^{o}\right\rangle\right\}$
@@
    (2) $\begin{split}\textbf{M}_{B}^{o}\rightarrow\textbf{M}_{B}^{s}=\left\{\left\langle x_{i}^{s},y_{i}^{s}\right\rangle\right\}_{i=1}^{l}&=\left\{\left\langle x_{1}^{s},y_{1}^{s}\right\rangle,\left\langle x_{2}^{s},y_{2}^{s}\right\rangle,\ldots,\left\langle x_{l}^{s},y_{l}^{s}\right\rangle\right\}\\ index_{R}&=I\left(S\left(\textbf{M}_{B}^{s}\right)\right)\end{split}$
@@
    (3) $\textbf{F}_{B}^{o}=\left\{\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{1}^{o}\right),y_{1}^{o}\right\rangle,\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{2}^{o}\right),y_{2}^{o}\right\rangle,\ldots,\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{l}^{o}\right),y_{l}^{o}\right\rangle\right\}$
@@
    (4) $\begin{split}\textbf{F}_{B}^{o}\rightarrow\textbf{F}_{B}^{s}&=R\left(\left\{\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{i}^{o}\right),y_{i}^{o}\right\rangle\right\}_{i=1}^{l},index_{R}\right)\\ &=\left\{\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{1}^{s}\right),y_{1}^{s}\right\rangle,\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{2}^{s}\right),y_{2}^{s}\right\rangle,\ldots,\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{l}^{s}\right),y_{l}^{s}\right\rangle\right\}\end{split}$
    Figure 2. The PDF of the U-shaped Beta probability distribution $\text{Beta}\left(\alpha,\alpha\right)$ corresponding to different shape parameters $\alpha$.
@@
    Table 1. The impact of different weight sampling numbers on AMPLIFY.
@@
  • Step 5: Perform element-wise Mixup operation on $\textbf{F}_{B}^{o}$ and $\textbf{F}_{B}^{s}$ to obtain the aggregated feature sequence set $\textbf{M}_{B}^{\text{mix}}$, which is then fed into the subsequent hidden neural network layer. Considering that AMPLIFY requires Mixup operation to be performed in each Transformer block, if the weight coefficients $\lambda$ for Mixup in different blocks are not the same, it will lead to frequent and intense disturbance to the features in the sequence, and some abnormal features or noise will also be constantly accumulated and strengthened, resulting in significant fluctuations in the model’s loss value during the actual experimental process. Moreover, when there is a large difference in the $\lambda$ values between each block, this instability will be further amplified. Therefore, the $\lambda_{\text{max}}$ we use is different from the definition in the standard Mixup. To improve the stability of linear interpolation and the consistency of feature representation, in the AMPLIFY algorithm, we adopt a method of first performing multiple samplings based on the Beta probability distribution (BPD), and then selecting the maximum value from the resulting $\lambda$ values. Specifically, we call function $\text{Beta}\left(\alpha,\alpha\right)$ multiple times to obtain a weight set with $n$ elements, denoted as $\Lambda$, and then select the maximum weight value $\lambda_{\text{max}}$ in this set as the weight coefficient for all Mixup operations in the model. This method can significantly reduce the adverse impact of randomness in the feature sequence on linear interpolation, making one feature sequence the explanatory term and the other the random perturbation term, which is also the reason why we tend to choose smaller $\alpha$ values. In addition, according to (Zhang et al., 2017)’s research, when $\alpha\rightarrow\infty$, the value of the BPD will approach 0.5 infinitely, and the training error of the model on the real dataset will also increase significantly. If the standard Mixup calculation method is used for mixing operation, it is likely to produce a large number of abnormal features due to significant disturbance caused by linear interpolation, which, in turn, will make the model prediction to deviate and greatly reduce its generalization performance. On the other hand, as shown in Figure 2, since the weight value obtained by randomly sampling only once from the U-shaped BPD may fall into the low probability area, or even be very close to 0.5. Therefore, sampling a total of $n$ weights from the BPD and taking the maximum value can effectively avoid this issue. According to experimental results, we also found that if the value of $n$ is large, the weight values sampled from the U-shaped BPD will appear in the high probability area in large quantities, and the maximum value selected from them will also be closer to 1. This results in a significant reduction in the weight of the feature sequence $\textbf{F}_{B}^{o}$ or $\textbf{F}_{B}^{s}$ that serves as the random perturbation term, and even renders it ineffective as a perturbation, thus making Mixup meaningless. Eventually, based on the experimental results 1 (We compared the classification performance of AMPLIFY with different weight sampling numbers on five datasets. The mean in the table is the average performance on the five datasets, and it reaches its optimal when sampled five times), we choose $a=0.1$ and $n=5$ to avoid obtaining risky weight values as much as possible and effectively leverage the role of random perturbation term. The above process can be described by the following equations:

### v3

@@ Abstract
Mixup is an effective data augmentation method that generates new augmented samples by aggregating linear combinations of different original samples. However, if there are noises or aberrant features in the original samples, Mixup may propagate them to the augmented samples, leading to over-sensitivity of the model to these outliers Visualization of Experimental Results. To solve this problem, this paper proposes a new Mixup method called AMPLIFY. This method uses the Attention mechanism of Transformer itself to reduce the influence of noises and aberrant values in the original samples on the prediction results, without increasing additional trainable parameters, and the computational cost is very low, thereby avoiding the problem of high resource consumption in common Mixup methods such as Sentence Mixup Guo et al., 2019a . The experimental results show that, under a smaller computational resource cost, AMPLIFY outperforms other Mixup methods in text classification tasks on 7 benchmark datasets, providing new ideas and new ways to further improve the performance of pre-trained models based on the Attention mechanism, such as BERT, ALBERT, RoBERTa, and GPT. Our code can be obtained at https://github.com/kiwi-lilo/AMPLIFY.
Introduction
@@
Data augmentation techniques are widely used in modern machine learning field to improve the predictive performance and robustness of computer vision and natural language processing (NLP) models by adding feature noise to the original samples. Compared to computer vision tasks, NLP tasks face samples with more complex data structures, semantic features, and semantic correlations, as natural language is the main channel of human communication and reflection of human thinking. Li et al., (2022) Therefore, NLP models are more sensitive to the quality of the dataset, and often require a variety of data augmentation techniques to improve the model’s generalization ability, adaptability, and robustness to new data in practical engineering applications. Common text data augmentation techniques include random synonym replacement  Zhang et al., (2015), back-translation  Xie et al., (2020), random word insertion and deletion, etc.  Wei and Zou, (2019). The core idea of these methods is to generate more training samples by performing a series of feature transformations on the original samples, allowing the model to learn and adapt to changing feature information, and enabling the model to better deal with and process uncertainty in the samples. Additionally, by selectively increasing the size of the training set, data augmentation techniques can effectively alleviate the negative impact of limited data and imbalanced data classification on model prediction performance.
@@
Standard Mixup is a simple and effective image augmentation method first proposed in 2017  Zhang et al., (2017). It aggregates two images and their corresponding labels by linear interpolation to generate a new augmented image and its pseudo-label. The main advantage of the standard Mixup technique is that it purposefully generates beneficial noise by using the weighted average of features in the original samples. After adapting to this noise, the model becomes less sensitive to other noise in the training samples, thereby improving the model’s generalization ability and robustness. Additionally, because it can generate more augmented samples by aggregating different original sample pairs, even at a high data augmentation magnitude, there is no duplicate aggregation result. This greatly enhances the diversity of augmented samples, thereby effectively improving the training efficiency of the model and reducing the risk of overfitting.In the NLP field, the standard Mixup technique can be broadly divided into two categories: Input Mixup  Yun et al., (2019) and Hidden Mixup  Verma et al., (2019). Input Mixup increases the number of samples in the training set input to the model. The specific process involves padding all text sequences in the training dataset to the same length. After completing word embedding processing, random pairs of sequences are combined. Mixup operations are then applied to the two vectorized samples in each sequence pair, along with their corresponding classification labels, generating augmented samples and their pseudo-labels. Finally, the augmented samples and pseudo-labels are fed into the model for training. Input Mixup performs Mixup operations during the word embedding stage of text preprocessing and only involves semantic features in the word vector space. Its main objective is to increase the training set size and reduce overfitting. Hidden Mixup applies Mixup operations in the hidden layers of the model. The specific process involves randomly selecting two sequences, $x$ and $x^{\prime}$, from the training set. After processing through the model, a random layer is selected from the model’s hidden layers, and the output features, $h$ and $h^{\prime}$, of $x$ and $x^{\prime}$ at that layer are extracted. Mixup is then performed on $h$, $h^{\prime}$, and their labels, resulting in linearly interpolated mixed features and corresponding pseudo-labels.
@@
Finally, the mixed features are fed into subsequent hidden layers. Hidden Mixup performs Mixup during the feature extraction stage of text processing and only involves specific hidden layers in the model. Its main objective is to enhance the model’s generalization ability without requiring additional input data. In summary, Input Mixup increases the sample quantity, while Hidden Mixup increases feature diversity. Both methods focus on improving model performance through Mixup operations, albeit with different emphases.  Verma et al., (2019)’s research has shown that performing Mixup operations at the input layer of the model leads to underfitting, while performing Mixup at deeper layers makes the training easier to converge. However, when the original samples contain noise or outliers, performing Mixup may result in generated augmented samples containing more regular noise or outliers due to the principle of linear interpolation. If the model learns too much from this noise and outliers, it may cause obvious overfitting and prediction bias, thereby reducing the model’s generalization performance. The basic idea of Standard Mixup is to linearly aggregate two samples $\left(x_{i},y_{i}\right)$ and $\left(x_{j},y_{j}\right)$ in the training set $\mathbb{D}_{train}$, to generate a new sample $\left(x_{\mathrm{mix}},y_{\mathrm{mix}}\right)$ as an augmented input to train the network, where $x$ represents the sample and $y$ represents its corresponding label. This aggregation process can be expressed by the following formulae:
@@
Figure 1: A schematic of AMPLIFY. In each encoder block of the Transformer, the forward-propagated input data is duplicated and re-ordered according to the label-mixing order after obtaining the results of the Muti-Head Attention, and then Mixup operation is performed. No changes are made to other network structures. Similarly, each decoder block can also perform the same operation.
@@
Multi-Head Attention (MHA) Vaswani et al., (2017) is a technique commonly used in sequence-to-sequence models to calculate feature correlations. Namely, it can calculate the attention weights for each position in the sequence in parallel using different attention heads, and then obtain the representation of the entire sequence by summing the weighted values. Specifically, MHA helps the model establish correspondences between features in the input and output sequences, thereby improving the expressiveness and accuracy of the model. It can also attend to different features at different positions in the input sequence without taking into account their distance, assigning different weights to different parts of the input to capture information relevant to the output sequence. To address the problem of noise and out-of-distribution values in original samples, which affect the aggregated samples in standard Mixup, our AMPLIFY method uses the Hidden Mixup idea to perform Mixup operations on the hidden layers of the Transformer block. Since the output of the MHA already includes the model’s attention to different parts of the input sequence, it can guide the model on which features to retain and which to discard when aggregating different samples, ultimately integrating more critical features into the generated output sequence. Therefore, AMPLIFY duplicates the MHA output results of the sample sequence in the same batch, shuffles the order of these copies, and then performs Mixup on them with the original output results. This allows the model to aggregate the correlations and attention of two sample features multiple times at different levels, thereby obtaining more reasonable weights for each position in the aggregated sequence and avoiding unnecessary noise or feature information loss during the Mixup process. Additionally, since the Attention mechanism itself is a method of weighting the input features, AMPLIFY is more natural and effective intuitively.The overall structure of the model is depicted in Figure 1
@@
Since AMPLIFY aggregates the features of the augmented sample and the corresponding original sample in the hidden layers of the model, it has lower computational cost and fewer parameters compared to other data augmentation methods. It also avoids the problem of increased time and resource consumption caused by standard Mixup methods. In addition, Our result have shown that our method can better learn the key features of the input sequence, improving the generalization ability and prediction performance of the model. For example, on the MRPC dataset, AMPLIFY’s average accuracy is 1.83% higher than the baseline, and 1.04%, 1.72%, and 1.46% higher than EmbedMix, SentenceMix, and TMix, respectively.
Motivation
@@
Data Augmentation: refers to a class of methods that generate new data with certain semantic relevance based on the features of existing data. By adding more augmented data, the overall prediction performance of the model can be improved and the robustness of the model can be enhanced. In the case of a limited size of the training set, data augmentation techniques are more effective because they can significantly reduce the risk of overfitting and improve the generalization ability of the model. However, the inherent difficulty of NLP tasks makes it difficult to construct data augmentation methods similar to those in the computer vision field (such as cropping and flipping), which may significantly alter the semantics of the text and make it difficult to balance the quality of the data with the diversity of the features. Currently, data augmentation methods in the NLP field can be roughly divided into three categories: Paraphrasing  Wang and Yang, (2015), Noising  Wei and Zou, (2019), and Sampling  Min et al., (2020). Among them, Paraphrasing methods involve certain transformations of the characters, words, phrases, and sentence structures in the text while trying to retain the semantics of the original text. However, such methods may lead to differences in textual semantics in different contexts, for example, substituting ”I eat an iPhone every day” for ”I eat an apple every day” obviously does not make sense. Noising methods aim to add some continuous or discrete noise to the sentence while keeping the label unchanged. Although this method has little impact on the semantics of the text, it may have a significant impact on the basic structure and grammar of the sentence, and there are also certain limitations in improving the diversity of features. For example, adding noise to the sentence ”i like watching sunrise” to turn it into ”i like, watching jogging sunrise” would impair the grammatical integrity of the original sentence. The goal of Sampling methods is to select samples based on the feature distribution of the existing data and use these samples to augment new data. This method needs to define different selection strategies manually according to the features of different datasets, so its application range is limited and the diversity of the features obtained by augmentation is relatively poor.
@@
Mixup: is a noise-based data augmentation technique introduced by  Zhang et al., (2017). It was first introduced in the field of computer vision and is now widely used in the entire field of deep learning. Its purpose is to improve the predictive performance of neural network models by generating regularly noised samples. The standard Mixup method generates a new augmented image by linearly interpolating two existing images and their labels, encouraging the model to learn more generalized decision boundaries. To bring the idea of standard Mixup data augmentation to the NLP field,  Guo et al., 2019a modified the underlying operations of standard Mixup and proposed two methods for applying Mixup in NLP tasks: EmbedMix, which performs Mixup in the word embedding space, and SentenceMixup, which performs Mixup after information aggregation in the final hidden layer.  Sun et al., (2020) first proposed a strategy for doing Mixup on transformer-based pre-trained text models and concluded that the smaller the data volume, the greater the improvement in model performance due to Mixup.  Chen et al., (2020) proposed a Mixup method for semi-supervised text classification called TMix, which uses two different sets of data to randomly select a layer in all hidden layers of the model and perform Mixup on its features. SeqMix, proposed by  Zhang et al., 2020b , is a technique for improving feature diversity during active learning by mixing sequences in the feature space of hidden layers. It provides an effective solution for Mixup at the subsequence level and a method for augmenting the labels of active sequences.
@@
Attention: is a commonly used technique in machine learning that allows models to focus on important information in sequence data and learn the underlying features of that information. In the NLP field, Attention assigns a weight to each hidden state and determines which feature state the output should focus on based on the varying weights. This helps the model better understand key features in a sentence, ultimately improving the performance of sequence models. Recently, some computer vision models have attempted to combine Mixup with saliency-based Attention mechanisms, such as attentive-CutMix  Walawalkar et al., (2020), puzzle-Mix  Kim et al., (2020), and saliency-CutMix  Uddin et al., (2020). One successful case of introducing Mixup to the mainstream pre-trained model, Vision Transformer (ViT), is TransMix  Chen et al., (2022), which dynamically reassigns label weights based on the response of each data point in the attention map. Inputs that are better focused in the attention map are assigned higher values in the mixed labels.
@@
Hidden Mixup: Recent research in data augmentation has explored the use of Mixup in the hidden layers of deep learning models. For example,  Verma et al., (2019) proposed a Mixup augmentation algorithm called Manifold Mixup, which trains neural networks on interpolated hidden representations and encourages the network to maintain uncertainty about the size of the unobserved feature representation space during training. This causes the feature representations of real training samples to be concentrated in a low-dimensional subspace, resulting in more discriminative aggregated features.  Yoon et al., (2021) proposed a data augmentation algorithm called SsMix, which performs span-based mixing on input text sequences. This method selects the least salient span in text $A$ and replaces it with the most salient span of the same length in text $B$, while preserving most of the important tokens in both sequences. However, current research has not yet explored the combination of Mixup and MHA in the hidden layer of Transformer architecture.
Method
Portions of this text were previously published as part of a preprint (https://arxiv.org/abs/2309.12689)
@@
    $\textbf{M}_{B}^{o}:\mathbb{D}_{train}\rightarrow\left\{\left\langle x_{i}^{o},y_{i}^{o}\right\rangle\right\}_{i=1}^{l}=\left\{\left\langle x_{1}^{o},y_{1}^{o}\right\rangle,\left\langle x_{2}^{o},y_{2}^{o}\right\rangle,\ldots,\left\langle x_{l}^{o},y_{l}^{o}\right\rangle\right\}$ (1)
@@
    $\begin{split}\textbf{M}_{B}^{o}\rightarrow\textbf{M}_{B}^{s}=\left\{\left\langle x_{i}^{s},y_{i}^{s}\right\rangle\right\}_{i=1}^{l}&=\left\{\left\langle x_{1}^{s},y_{1}^{s}\right\rangle,\left\langle x_{2}^{s},y_{2}^{s}\right\rangle,\ldots,\left\langle x_{l}^{s},y_{l}^{s}\right\rangle\right\}\\ index_{R}&=I\left(S\left(\textbf{M}_{B}^{s}\right)\right)\end{split}$ (2)
@@
    $\textbf{F}_{B}^{o}=\left\{\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{1}^{o}\right),y_{1}^{o}\right\rangle,\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{2}^{o}\right),y_{2}^{o}\right\rangle,\ldots,\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{l}^{o}\right),y_{l}^{o}\right\rangle\right\}$ (3)
@@
    $\begin{split}\textbf{F}_{B}^{o}\rightarrow\textbf{F}_{B}^{s}&=R\left(\left\{\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{i}^{o}\right),y_{i}^{o}\right\rangle\right\}_{i=1}^{l},index_{R}\right)\\ &=\left\{\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{1}^{s}\right),y_{1}^{s}\right\rangle,\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{2}^{s}\right),y_{2}^{s}\right\rangle,\ldots,\left\langle A_{\text{pre}}^{\text{MHA}}\left(x_{l}^{s}\right),y_{l}^{s}\right\rangle\right\}\end{split}$ (4)
    Figure 2: The PDF of the U-shaped Beta probability distribution $\text{Beta}\left(\alpha,\alpha\right)$ corresponding to different shape parameters $\alpha$.
@@
    Table 1: The impact of different weight sampling numbers on AMPLIFY.
@@
  • Step 5: Perform element-wise Mixup operation on $\textbf{F}_{B}^{o}$ and $\textbf{F}_{B}^{s}$ to obtain the aggregated feature sequence set $\textbf{M}_{B}^{\text{mix}}$, which is then fed into the subsequent hidden neural network layer. Considering that AMPLIFY requires Mixup operation to be performed in each Transformer block, if the weight coefficients $\lambda$ for Mixup in different blocks are not the same, it will lead to frequent and intense disturbance to the features in the sequence, and some abnormal features or noise will also be constantly accumulated and strengthened, resulting in significant fluctuations in the model’s loss value during the actual experimental process. Moreover, when there is a large difference in the $\lambda$ values between each block, this instability will be further amplified. Therefore, the $\lambda_{\text{max}}$ we use is different from the definition in the standard Mixup. To improve the stability of linear interpolation and the consistency of feature representation, in the AMPLIFY algorithm, we adopt a method of first performing multiple samplings based on the Beta probability distribution (BPD), and then selecting the maximum value from the resulting $\lambda$ values. Specifically, we call function $\text{Beta}\left(\alpha,\alpha\right)$ multiple times to obtain a weight set with $n$ elements, denoted as $\Lambda$, and then select the maximum weight value $\lambda_{\text{max}}$ in this set as the weight coefficient for all Mixup operations in the model. This method can significantly reduce the adverse impact of randomness in the feature sequence on linear interpolation, making one feature sequence the explanatory term and the other the random perturbation term, which is also the reason why we tend to choose smaller $\alpha$ values. In addition, according to Zhang et al., (2017)’s research, when $\alpha\rightarrow\infty$, the value of the BPD will approach 0.5 infinitely, and the training error of the model on the real dataset will also increase significantly. If the standard Mixup calculation method is used for mixing operation, it is likely to produce a large number of abnormal features due to significant disturbance caused by linear interpolation, which, in turn, will make the model prediction to deviate and greatly reduce its generalization performance. On the other hand, as shown in Figure 2, since the weight value obtained by randomly sampling only once from the U-shaped BPD may fall into the low probability area, or even be very close to 0.5. Therefore, sampling a total of $n$ weights from the BPD and taking the maximum value can effectively avoid this issue. According to experimental results, we also found that if the value of $n$ is large, the weight values sampled from the U-shaped BPD will appear in the high probability area in large quantities, and the maximum value selected from them will also be closer to 1. This results in a significant reduction in the weight of the feature sequence $\textbf{F}_{B}^{o}$ or $\textbf{F}_{B}^{s}$ that serves as the random perturbation term, and even renders it ineffective as a perturbation, thus making Mixup meaningless. Eventually, based on the experimental results Table 1 (We compared the classification performance of AMPLIFY with different weight sampling numbers on five datasets. The mean in the table is the average performance on the five datasets, and it reaches its optimal when sampled five times), we choose $a=0.1$ and $n=5$ to avoid obtaining risky weight values as much as possible and effectively leverage the role of random perturbation term. The above process can be described by the following equations: